OpenStudy (anonymous):

Leah purchased a lawn mower for $3,865. It depreciates 1.5% each year. What is the value of the lawn mower after seven years?

OpenStudy (anonymous):

This is a classic decay problem: \[y = initial * (1-decay)^(n-1)\] so for you, the equation is \[value=3865 * (.985)^(year-1)\] This way, each time you add in a year, the intitial value is multiplied by an additional .985, effectively lowering the price by 1.5%. The -1 is important because it makes it so at year 1, there is no decay yet, not until year 2. Year 1 is 3865 * .9865^0 = 3865 * 1 = 3865, which makes sense because it's your initial value ;) So anyways, to find at 7 years just plug in 7 for your year :)

OpenStudy (anonymous):

Those are supposed to be raised to the n-1 and raised to the year-1, not sure what happened there, lol

OpenStudy (anonymous):

i got 3529

OpenStudy (anonymous):

wht did u get

OpenStudy (anonymous):

Actually, thinking about it now, your question is worded as "after _ years", so you don't want the -1 there. Your answer would be right otherwise. After 1 year, you do want decay, I was thinking at year 1 :) So just multiply your answer by 1 more .985 to get the real answer